two men P and Q are standing at corners A and B of square ABCD of side 8m. they start moving along the track with constant speed 2m/s and 10 m/s respectively./ find the time after which they will meet

Grade:12th pass

4 Answers

Anoopam Mishra
8 years ago
P and Q are walking at speed of 2m/s and 10m/s.
This situation is similar to that when P is at rest and Q walks at 8m/s. Q has to walk 3*8 = 24 m for which he will take 24/8 = 3 seconds.
Hence P and Q will meet after 3 seconds.
mukul agrawal
8 years ago
how you calculate the distance =24. we have only speed and no time,then?
 
Anoopam Mishra
8 years ago
P and Q are at the A and B. We have assumed that P is at rest
Anoopam Mishra
8 years ago
P and Q are at the A and B. We have assumed that P is at rest and Q walks at 8m/s. Now Q has to walk from B to C, from C to D, from D to A to reach P. We needs to cover a distance equal to 3 times the length of one side of the square. Hence distance = 8*3 = 24 m.
If you have any doubts please reply.
Please approve.